OpenStudy (anonymous):

The sum of the length, width, and height of a rectangular box is 16 cm; the width is twice the height; and twice the length exceeds the sum of the width and height by 5. Find the length, width, and height of the box using matrices. A. Length = 10 cm, Width = 4 cm, Height = 2 cm B. Length = 7 cm, Width = 6 cm, Height = 3 cm C. Length = 7 cm, Width = 3 cm, Height = 6 cm D. Length = 10 cm, Width = 2 cm, Height = 4 cm

OpenStudy (anonymous):

l+w+h=16 w=2h 2l=5+w+h (5+2h+h)/2+2h+h=16 5/2+3h/2+3h=16 5+3h+6h=32 9h=27 h=3 w=6 l=7

OpenStudy (zale101):

Let us not help the users by giving out the explanation and then the direct answer. Step by step is the effective process when tutoring. Let the tutor hold the pen and you guide them through is what they told me when i tutor at my school.